OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x). @jordanloveangel

OpenStudy (jordanloveangel):

assuming that x2= x^2 and (f-1)(x) f(x) = x^2 -16 (f-1)(x) = x^2-x-16

OpenStudy (jordanloveangel):

do u understand it now

OpenStudy (anonymous):

yeah that makes more since. Thanks, Mind helping me with a few more?

OpenStudy (jordanloveangel):

sure